OpenStudy (anonymous):

How do you find the x-intercept of f(x)=x^2-7 and f(x)=(x+4)^2-3?

OpenStudy (anonymous):

the x-intercept is when y=0 so you need to set each equation equal to 0 and solve.

OpenStudy (anonymous):

Do you know how to do this?

OpenStudy (anonymous):

do i just solve for x or do i need to use the quadratic formula?? but is that possible since there's no b

OpenStudy (anonymous):

Lets work through it. Take the left equation first. If you set it equal to 0 and solve for x we get to \[7=x ^{2}\]

OpenStudy (anonymous):

okay then do i square root the 7?

OpenStudy (anonymous):

from here we can say that \[x=\pm \sqrt{7}\] because when a negative is squared it becomes a positive, so there are two answers for the first equation. write this in coordinate form.

OpenStudy (anonymous):

okay then. then it's the same for the other function too?

OpenStudy (anonymous):

its a little trickier but ya. set it equal to zero and start to solve for x. you will have to add 3 to each side then expand the paranthesis

OpenStudy (anonymous):

nevermind about expanding that will make it more complicated

OpenStudy (anonymous):

oh okay then. should i expand the parentheses first then add 3 to both side??

OpenStudy (anonymous):

haha oh okay then then just add 3, square toot it, then - 4?

OpenStudy (anonymous):

hold on. I graphed it now I need to prove it arithmatically

OpenStudy (anonymous):

haha now im confused

OpenStudy (anonymous):

there needs to be 2 answers . x = -2.268, -5.732 by calculator. I'm sorry this one has me stumped about solving it by hand,

OpenStudy (anonymous):

the first is correct

OpenStudy (anonymous):

haha i actually have the answer in the back of the book but i dont get how to do it

OpenStudy (anonymous):

I think you can say x+4 needs to equal the sqrare root of 3

OpenStudy (anonymous):

@ChmE is right.

OpenStudy (anonymous):

and solve it for x+3 = - square root of 3.

OpenStudy (anonymous):

ahh scary perso here >.>.....

OpenStudy (anonymous):

this should give you both of your answers

OpenStudy (anonymous):

you suppose to set the y to zero and solve for the x. haha whatever @Cookiencreams17 and he best way to do it is reverse pemdas.

OpenStudy (anonymous):

you have to solve for both positive and negative because when it is squared it wont matter. This is the correct method.

OpenStudy (anonymous):

yep it's correct but the're 2 points

OpenStudy (anonymous):

oh cool okay i got it thanks for the help

OpenStudy (anonymous):

Ya, That one tripped me up for a bit. lol. welcome

OpenStudy (anonymous):

if \[3=(x+4)^{2}\]then we can say \[+\sqrt{3}=x+4\]and\[-\sqrt{3}=x+4\]because when you take the square of each we arrive back at the original equation. If you solve for both of these equations you should get the 2 answers for x

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yea it heps a lot thank you again. oh i was wondering how you do the square root thing haha i think it's cool

OpenStudy (anonymous):

what thing?

OpenStudy (anonymous):

the square root symbol

OpenStudy (anonymous):

oh, click on the equation button below this text box. you can do all sorts of things

OpenStudy (anonymous):

oh thats cool